OpenStudy (anonymous):

(square root of 7 - square root of 11)^2

OpenStudy (kaederfds):

18 - 2 square root of 77

OpenStudy (ash2326):

\[(\sqrt 7 -\sqrt {11})^2\] Use the formula \[(a-b)^2=a^2-2ab+b^2\] \[(\sqrt 7 -\sqrt {11})^2=(\sqrt 7)^2- 2\times \sqrt 7 \sqrt{11} +(\sqrt {11})^2\] \[(\sqrt 7 -\sqrt {11})^2=7-2 \times \sqrt {77}+11\] Can you simplify now?

OpenStudy (anonymous):

16*\[\sqrt{11}\]

OpenStudy (ash2326):

@ngtts \[ 7 -( 2\times \sqrt {11} )+11\] Now simplify it:)